10) Suppose that the effective return to a U.S. investor from buying a U.K. bond is 5.55%. Forward and spot exchange rates ($/£) are 2.10 and 2.00 respectively. The interest rate on the U.K. bond is most likely equal to: A) 5.45% B) 5.500% C) 5.650% D) 5.60%
Answer #1

Answer is option B)

from interest rate parity condition

i = i* + ∆e/e

i* = UK interest rate

i = effective return to US investor, investing in Pounds

∆e/e = expected appreciation of Pound

∆e/e = (FR-SR)/SR

= (2.1-2)/2 = .05

as over the time, Pound appreciates

so,

5.55 = i* + .05

i* = 5.55-.05

= 5.50

so i in UK = 5.500%